OpenStudy (anonymous):

The equation of a line is y = 2x – 1. What is the equation of the line that is parallel to the first line and passes through (–4, 1)? Anyone help please

OpenStudy (mrnood):

OK Do you understand the form of the equation y= mx + b? What is the slope of the line you are given?

OpenStudy (anonymous):

is it 2?

OpenStudy (mrnood):

yes and all parallel lines have the SAME slope the the line you want to find is y= 2x + b you need to find b but the point (-4,1) is on the line so 1= 2(-4) +b so oyu can find b

OpenStudy (anonymous):

So do i just solve for 1=2(-4)+b?

OpenStudy (mrnood):

yes- that will give oyu the value of b and you already know that m = 2 fro the line so the answer is y=2x+b

OpenStudy (anonymous):

A. y = 2x – 3 B. y = –2x – 7 C. y= -1/2x-1 D. y = 2x + 9

OpenStudy (anonymous):

those are the possible answers that i have

OpenStudy (mrnood):

OK - so solve the equation for b and you have it

OpenStudy (anonymous):

B=-2x+y?